Friday 08/03/12

9:50 A – 8:15 P

  • STORM O/N run still scanning (ran 488 movie 2Hz too long).  Aborted
  • Make new Glox (STORM4 images look dim / not enough photons/switches given brightness of conventional signal).  suspect Glox kept in door might have undergone too much freeze thaw, people like to leave that freezer door open for ages.
  • Making new sparser 100nm bead slide for imaging. First attempt aspirate beads and wash with water leads to beads not stuck to surface.  Repeating with brief wash off and air dry technique.
  • KE in situs day 2
  • saving Zeiss Confocal data as 12 bit doesn’t seem to reduce image size, still exported as 16bit.
  • Processing m197 and m198 data from Zeiss700 imaging.
  • centromere K27 Dm01 stains look great on Turnkey.
  • Sebastian practice talk 3p-4:30p
  • S5 sample substantially under-etched at 22min
  • Running v3 of promoter snail simulation — explicit time to turn on per cell, gamma random variable. ON once is ON for cell cycle (still needs pol II binding and enhancer activation).
  • Etch S2 for 45 min — still see resin.  etch another 20 min and mount.  8:10P.
  • centromere K27 Dm01 stain post-fixed and now in ETOH at 4C
  • STORM4 off.
This entry was posted in Summaries and tagged , . Bookmark the permalink.